# MSACL: Multi-Step Actor-Critic Learning with Lyapunov Certificates for Exponentially Stabilizing Control

### v1

@@ Abstract
Achieving provable stability guarantees in model-free reinforcement learning (RL) for control systems remains a fundamental challenge, particularly in balancing exploration efficiency with rigorous safety requirements. This article introduces MSACL, a novel framework that seamlessly integrates exponential stability theory with maximum entropy reinforcement learning through multi-step Lyapunov certificate learning. Unlike conventional approaches that rely on complex reward engineering or single-step constraints, MSACL leverages off-policy multi-step trajectory data to learn Lyapunov certificates that strictly satisfy theoretical stability conditions. By introducing the concept of Exponential Stability Labels (ESL) and a $\lambda$-weighted aggregation mechanism, the framework effectively balances the bias-variance trade-off inherent in multi-step learning. Policy optimization is subsequently guided by a stability-aware advantage function, which ensures that the learned policy $\pi$ promotes the rapid descent of the Lyapunov certificate. We rigorously evaluate the proposed method across six benchmark environments, encompassing four stabilization tasks and two high-dimensional nonlinear tracking scenarios. The experimental results demonstrate consistent superiority over baseline RL algorithms and state-of-the-art Lyapunov-based RL algorithms. Remarkably, MSACL achieves both exponential stability and rapid convergence under simple, intuitive reward structures, while exhibiting significant robustness to environmental uncertainties and exceptional generalization to unseen reference trajectories. Our sensitivity analysis reveals that the multi-step horizon $n$ is a critical factor in stability learning, with $n=20$ established as a robust default across diverse dynamical systems. By establishing a rigorous connection between Lyapunov stability theory and off-policy actor-critic frameworks, MSACL provides a foundational approach for developing verifiably safe learning-based controllers in safety-critical applications. To facilitate reproducibility and further research, the source code and benchmark environments of MSACL will be made publicly available soon.
@@ I Introduction
Reinforcement learning (RL) has emerged as a powerful paradigm for controlling complex, nonlinear dynamical systems without requiring analytical models [1]. Despite its empirical success in high-dimensional tasks [2, 3], the widespread deployment of RL in safety-critical applications, including autonomous driving and human-robot interaction, remains constrained by the lack of formal reliability guarantees [4, 5]. In these domains, stability is not merely a desirable metric but an imperative constraint; failure to maintain the system within a safe operating region can lead to catastrophic consequences. While approaches like constrained Markov Decision Processes (cMDPs) [6, 7] and Hamilton-Jacobi reachability analysis [8, 9] offer partial solutions, certificate-based methods have established themselves as the gold standard for verifying learning agents. By leveraging rigorous control-theoretic tools, specifically Lyapunov functions for asymptotic stability [10, 11] and control barrier functions (CBFs) for set invariance [12, 13], these methods provide concise, data-driven proofs that guarantee the system will converge to an equilibrium or remain strictly within safety boundaries, offering a level of assurance that standard reward-driven learning cannot match.
The synthesis of stability and safety certificates has undergone a paradigm shift from computationally expensive analytical methods, such as sum-of-squares (SOS) optimization with its inherent restriction to polynomial dynamics [14], to scalable neural approaches. Around 2018, foundational work on neural certificates emerged, marking a turning point in bridging black-box learning with verifiable control. Early efforts by Berkenkamp et al. [15] employed Gaussian Processes (GP) to learn Lyapunov functions for safe exploration, while Richards et al. [16] and Chow et al. [17] developed theoretical frameworks for synthesizing Lyapunov functions to estimate regions of attraction. These studies established that Lyapunov certificates as well as barrier-based counterparts provide a rigorous mathematical foundation for analyzing the stability and safety of nonlinear systems. Building on this verification-centric perspective, recent research has advanced toward the joint synthesis of policies and certificates: Chang et al. [18], for instance, showed that a neural certificate can act as a “critic” or supervision signal during training, directly guiding policy learning to satisfy formal stability conditions. This evolution underscores the dual role of learned certificates, which not only enable post-hoc safety verification but also actively shape the learning process to ensure stability and safety by design, offering guarantees that standard trial-and-error reinforcement learning cannot provide.
To enable effective learning in complex continuous spaces, we build our approach upon the framework of Maximum Entropy Reinforcement Learning (MERL). Algorithms such as Soft Actor-Critic (SAC) [19] and its distributional variants like DSAC [20, 21] maximize not just the expected return but also the entropy of the policy. This entropy regularization encourages robust exploration and prevents the policy from converging prematurely to suboptimal deterministic behaviors. Given its superior sample efficiency and robustness to disturbances, the maximum entropy framework serves as an ideal substrate for developing advanced, stability-aware control algorithms.
However, standard model-free RL algorithms, including DDPG [22], TRPO [23], PPO [24], TD3 [25], D4PG [26], SAC, and DSAC, face significant limitations when applied to stabilization or tracking tasks. First, they lack intrinsic stability guarantees; since policies trained via these methods optimize solely for scalar reward maximization, stability is essentially an incidental byproduct that is achieved only when the agent implicitly discovers stable behaviors during the exploration process, rather than being structurally enforced. Second, the efficacy of these algorithms relies heavily on exhaustive reward engineering, requiring practitioners to laboriously calibrate coefficients for state deviation, control effort, and auxiliary terms to elicit stable behavior. Our proposed framework, MSACL (Multi-Step Actor-Critic Learning with Lyapunov Certificates), directly addresses these shortcomings. By adopting a formulation rooted in quadratic performance functionals such as LQR-like costs, MSACL eliminates the need for heuristic reward shaping. Most importantly, it guarantees stability by design, ensuring that the synthesized controller is not only high-performing but also mathematically certifiable.
While Lyapunov-based RL algorithms have been proposed to mitigate the two issues above, existing methods exhibit distinct shortcomings. The Policy Optimization with Lyapunov Certificates (POLYC) [27] algorithm offers theoretical guarantees but relies on on-policy learning, which results in prohibitive sample complexity for physical systems. Conversely, the Lyapunov-based Actor-Critic (LAC) [28] algorithm introduces an off-policy approach but fundamentally misaligns the objective: it minimizes a stability risk rather than a comprehensive control cost. This often leads to aggressive controllers that stabilize the system at the cost of excessive control effort, rendering them impractical for real-world actuators. Moreover, POLYC and LAC typically certify only asymptotic stability, a relatively weaker stability notion that ensures convergence but lacks the guaranteed exponential rate of decay. In practice, this guarantees eventual convergence but fails to impose a minimum convergence rate, often resulting in sluggish transient responses unacceptable for high-performance tasks. To overcome these limitations, MSACL is built upon an off-policy framework [29, 30], significantly improving data efficiency by leveraging historical agent-environment interaction data. Unlike LAC, MSACL optimizes a comprehensive objective covering both state regulation and control effort while enforcing exponential stability, which guarantees faster convergence rates and rigorous stability margins.
Another critical innovation of our approach is the utilization of multi-step data. Existing RL and Lyapunov-based methods typically rely on single-step transitions to enforce learning constraints. This pointwise reliance is susceptible to significant discretization errors when approximating the continuous-time Lie derivative of the Lyapunov function, particularly in the presence of local noise. To mitigate this and capture long-horizon dynamics, we introduce a multi-step learning mechanism. By training on $n$-step trajectories, we effectively bridge the gap between temporal difference learning and rigorous stability verification. We integrate this multi-step data with exponential stability theory to form a robust constraint mechanism. This allows the agent to learn a Lyapunov function that certifies exponential convergence over trajectory segments, rather than just instantaneous descent, thereby reconciling rigorous stability requirements with the need for flexible exploration.
@@
In summary, this article presents a comprehensive RL framework for model-free exponentially stabilizing control. The main contributions are summarized as follows:
@@
    Exponential Stability Guarantee via Multi-Step Learning: We introduce the MSACL framework, which integrates exponential stability conditions into Lyapunov certificate learning using a multi-step data utilization mechanism. By employing Exponential Stability Labels (ESL), our approach enforces that the system state converges to the equilibrium exponentially, providing stronger performance guarantees than conventional asymptotic convergence methods.
@@
    Sample-Efficient Off-Policy Framework: Developed within the maximum entropy RL paradigm, our framework enables efficient off-policy learning. By incorporating trajectory-wise importance sampling to correct for distribution shifts, MSACL allows for effective reuse of historical interaction data, significantly reducing sample complexity compared to existing on-policy Lyapunov-based RL algorithms.
@@
    Comprehensive Benchmarking System: We developed a diverse suite of six nonlinear dynamical environments, spanning four stabilization tasks (e.g., VanderPol) and two high-dimensional tracking tasks (e.g., Quadrotor Trajectory Tracking), to rigorously evaluate and provide a standard for Lyapunov-based RL algorithms.
  4. (iv)
    Robustness and Generalization Verification: Our methodology demonstrates significant resilience to stochastic process noise and parametric uncertainties. Furthermore, in tracking scenarios, the learned neural policies exhibit robust generalization capabilities, maintaining high-precision performance even when navigating geometrically distinct, unseen reference trajectories.
  5. (v)
    Bias-Variance Trade-off Analysis: We provide a systematic analysis of the multi-step horizon parameter $n$ and its impact on stability learning. Through the introduction of $\lambda$-weighted aggregation, we effectively manage the trade-off between estimation bias (associated with small $n$) and sampling variance (associated with large $n$), establishing as a robust heuristic for diverse control applications.
The remainder of this article is organized as follows. Section II formulates the stabilization and tracking problems and establishes the theoretical foundations, including the criteria for exponential stability and the maximum entropy reinforcement learning framework. Section III details the methodology for learning the Lyapunov certificate, specifically describing the multi-step data collection strategy and the loss functions designed to enforce formal stability conditions. Section IV presents the learning process for the neural control policy, deriving the optimization objectives for the soft Q-networks, the policy network, and the temperature coefficient. Section V reports the experimental results, providing a comprehensive comparison with baseline algorithms across six benchmark environments, followed by an evaluation of robustness, generalization capabilities, and parameter sensitivity. Finally, Section VI provides a discussion on the underlying design philosophy and concludes the article.
@@ II Preliminaries
In this section, we establish the theoretical foundations for the proposed MSACL framework. We first define the stabilization and tracking problems for unknown discrete-time systems. Subsequently, the concept of exponential stability is introduced along with a Lyapunov-based certificate lemma, providing the criteria for convergence guarantees. Finally, we present the maximum entropy reinforcement learning framework, bridging Lyapunov-based control with model-free actor-critic optimization.
@@ II-A Problem Statement
In this article, we study the stabilization and tracking problems for general discrete-time deterministic control systems:
@@
where $\mathbf{x}_{t}\in\mathcal{X}\subseteq\mathbb{R}^{n}$ denotes the state at time $t$, $\mathbf{u}_{t}\in\mathcal{U}\subseteq\mathbb{R}^{m}$ represents the control input, $\mathcal{X}$ and $\mathcal{U}$ denote the sets of admissible states and control inputs, respectively, and $f:\mathcal{X}\times\mathcal{U}\to\mathcal{X}$ is the flow map. We assume that $f$ is unknown in our setting, thus focusing on model-free control. It is worth noting that the discrete-time framework presented in this article can be similarly extended to the continuous-time case; interested readers may refer to [10] and [27] for further details.
In the context of control systems, our objective is to find a state feedback policy $\pi:\mathcal{X}\to\mathcal{U}$ such that the control input $\mathbf{u}_{t}=\pi(\mathbf{x}_{t})$ renders the closed-loop system $\mathbf{x}_{t+1}=f(\mathbf{x}_{t},\pi(\mathbf{x}_{t}))$ to satisfy desired properties (e.g., stability, safety). For clarity, we consider the scenario where the state $\mathbf{x}_{t}$ is fully observable, and the policy $\pi$ is a time-invariant mapping. Specifically, we focus on tasks including
  • Stabilization: The objective is to find a policy $\pi$ such that the state ultimately converges to a target state $\mathbf{x}_{g}\in\mathcal{X}$. Without loss of generality, we assume $\mathbf{x}_{g}$ to be the origin, i.e., $\mathbf{x}_{g}=\mathbf{0}$.
  • Tracking: The goal is to find a policy $\pi$ such that the state $\mathbf{x}_{t}$ eventually converges to a reference signal $\mathbf{x}_{t}^{\text{ref}}$ (which may represent a desired trajectory, velocity, or attitude). We define the tracking error as $\mathbf{e}_{t}=\mathbf{x}_{t}-\mathbf{x}_{t}^{\text{ref}}$, thereby transforming the tracking task into a stabilization task where $\mathbf{e}_{t}$ must converge to zero.
In brief, the core problem addressed in this article is formulated as follows:
@@
For the control system (1), find a policy $\pi$ such that: (i) in stabilization tasks, $\mathbf{x}_{t}\to\mathbf{0}$; (ii) in tracking tasks, $\mathbf{e}_{t}\to\mathbf{0}$. Moreover, the policy should be optimized to achieve the fastest possible convergence rate (e.g., exponential stability).
@@ II-B Exponential Stability and Lyapunov Certificates
For the control system (1), the equilibrium point $\mathbf{x}_{g}=\mathbf{0}$ is said to be exponentially stable if there exist constants $C>0$ and $0<\eta<1$ such that
$\|\mathbf{x}_{t}-\mathbf{x}_{g}\|\leq C\eta^{t}\|\mathbf{x}_{0}-\mathbf{x}_{g}\|,\quad\forall\mathbf{x}_{0}\in\mathcal{X}_{0},$ (2)
where $\mathbf{x}_{0}\in\mathcal{X}_{0}\subseteq\mathcal{X}$ denotes the initial state and $t\in\mathbb{Z}_{\geq 0}$.
To guarantee that a closed-loop system is exponentially stable with respect to $\mathbf{x}_{g}$, we introduce the following lemma:
@@
for all $t\in\mathbb{Z}_{\geq 0}$, where $\mathbf{x}_{t+1}=f(\mathbf{x}_{t},\mathbf{u}_{t})$, then the origin $\mathbf{x}_{g}=\mathbf{0}$ is exponentially stable. The function $V$ is referred to as an exponentially stabilizing control Lyapunov function (ESCLF) [10, 31].
Proof: From the condition $V(\mathbf{x}_{t+1})\leq(1-\alpha_{3})V(\mathbf{x}_{t})$, we can iteratively derive the multi-step relationship:
$V(\mathbf{x}_{t+1})$ $\leq(1-\alpha_{3})V(\mathbf{x}_{t}),$ (4)
$V(\mathbf{x}_{t+2})$ $\leq(1-\alpha_{3})V(\mathbf{x}_{t+1})\leq(1-\alpha_{3})^{2}V(\mathbf{x}_{t}),$
$\vdots$
$V(\mathbf{x}_{t+n})$ $\leq(1-\alpha_{3})V(\mathbf{x}_{t+n-1})\leq\dots\leq(1-\alpha_{3})^{n}V(\mathbf{x}_{t}).$
By incorporating the lower and upper quadratic bounds $\alpha_{1}\|\mathbf{x}_{t}\|^{2}\leq V(\mathbf{x}_{t})\leq\alpha_{2}\|\mathbf{x}_{t}\|^{2}$, it follows that:
@@
Specifically, by setting the initial time to $t=0$ and denoting the current time step as $n=t\in\mathbb{Z}_{\geq 0}$, we obtain:
@@
Defining $C=\sqrt{\alpha_{2}/\alpha_{1}}$ and $\eta=\sqrt{1-\alpha_{3}}$, the result satisfies Definition 1, thereby proving that the control system (1) is exponentially stable with respect to $\mathbf{x}_{g}=\mathbf{0}$.

### v2

@@ Abstract
For safety-critical applications, model-free reinforcement learning (RL) faces numerous challenges, particularly the difficulty of establishing verifiable stability guarantees while maintaining high exploration efficiency. To address these challenges, we present Multi-Step Actor-Critic Learning with Lyapunov Certificates (MSACL), a novel approach that seamlessly integrates exponential stability with maximum entropy reinforcement learning (MERL). In contrast to existing methods that rely on complex reward engineering and single-step constraints, MSACL utilizes intuitive rewards and multi-step data for actor-critic learning. Specifically, we first introduce Exponential Stability Labels (ESLs) to categorize samples and propose a $\lambda$-weighted aggregation mechanism to learn Lyapunov certificates. Leveraging these certificates, we then develop a stability-aware advantage function to guide policy optimization, thereby ensuring rapid Lyapunov descent and robust state convergence. We evaluate MSACL across six benchmarks, comprising four stabilization and two high-dimensional tracking tasks. Experimental results demonstrate its consistent superiority over both standard RL baselines and state-of-the-art Lyapunov-based RL algorithms. Beyond rapid convergence, MSACL exhibits significant robustness against environmental uncertainties and remarkable generalization to unseen reference signals. The source code and benchmarking environments are available at https://github.com/YuanZhe-Xing/MSACL.
@@ I Introduction
Reinforcement learning (RL) has emerged as a transformative paradigm for the control of complex, nonlinear dynamical systems [1, 2], demonstrating remarkable empirical success across high-dimensional tasks without requiring explicit analytical models [3, 4]. Despite these advancements, the deployment of RL in safety-critical applications, such as autonomous driving and human-robot interaction, remains fundamentally constrained by a lack of formal reliability guarantees [5, 6].
While alternative frameworks like constrained Markov Decision Processes (cMDPs) [7, 8] and Hamilton-Jacobi reachability analysis [9, 10] offer potential solutions, certificate-based methods constitute a fundamental framework for verifying and stabilizing learning agents. By leveraging rigorous control-theoretic tools, such as Lyapunov certificates for establishing asymptotic stability [11, 12, 14, 13] and barrier certificates for ensuring set invariance [15, 16], these methods provide concise, data-driven proofs of system behavior. The synthesis of stability certificates has undergone a paradigm shift from computationally intensive analytical methods toward scalable neural architectures. Traditional approaches, such as sum-of-squares (SOS) optimization [17, 18, 19] and algebraic synthesis[20], are often restricted to polynomial dynamics. To bridge the gap between black-box learning and verifiable control, the field has evolved from probabilistic methods, notably the use of Gaussian Processes for safe exploration [21], to formal frameworks for synthesizing neural certificates that estimate regions of attraction (ROA) [22, 23]. Building on this verification-centric perspective, recent research has advanced toward the joint synthesis of neural certificates and policies. For instance, Chang et al. [24, 25] and Quan et al. [26, 27, 28] demonstrated that neural certificates can function as critics during training, directly guiding policy optimization to satisfy formal stability constraints.
In contrast to the stability-by-design paradigms presented above, standard model-free RL algorithms, such as DDPG [29], PPO [31], TD3 [32], SAC [33], and SAC’s distributional variant DSAC [34, 35], lack intrinsic stability guarantees, thus facing fundamental limitations in precision control tasks. These methods treat stability as an incidental byproduct of reward maximization rather than a structural property. Consequently, employing general RL algorithms for safety-critical tasks necessitates extensive reward engineering to achieve stable behaviors [1, 2]. To bridge this gap, Lyapunov-based RL algorithms have been developed to integrate stability constraints into the learning process. For instance, Policy Optimization with Lyapunov Certificates (POLYC) [25] incorporates Lyapunov certificates within the PPO framework, while Lyapunov-based Actor-Critic (LAC) [36] introduces an off-policy approach for learning stability-constrained critics. By augmenting standard RL architectures with learned Lyapunov certificates, these methods provide structural stability guarantees for nonlinear control systems. However, several critical limitations persist. First, these approaches suffer from low sample efficiency under on-policy training paradigms. Second, they often fail to account for control costs during policy optimization. More critically, most existing methods focus solely on asymptotic stability, which lacks the rapid stabilization characteristics necessary for high-performance response in dynamic control scenarios.
To overcome these limitations, we propose Multi-Step Actor-Critic Learning with Lyapunov Certificates (MSACL), a Lyapunov-based off-policy RL algorithm that enhances sample efficiency [37, 38] while optimizing a comprehensive objective that balances state regulation and control cost. By integrating exponential stability constraints into the maximum entropy reinforcement learning (MERL) architecture, MSACL ensures rapid stabilization and enhanced robustness. A key feature of our approach is the multi-step learning mechanism, which captures long-horizon dynamics to establish connections between temporal difference learning and formal stability verification. By mitigating discretization errors and stochastic noise associated with single-step transitions, this mechanism enables the synthesis of policies that are both high-performing and mathematically certifiable.
@@
In summary, this article presents a Lyapunov-based RL approach for model-free exponentially stabilizing control. The main contributions are summarized as follows:
@@
    Exponential Stability via Multi-Step Learning. We propose the MSACL algorithm, which integrates exponential stability with Lyapunov certificate learning via a multi-step mechanism. This algorithm enforces exponential convergence rates, providing stability-guaranteed performance for complex nonlinear control, surpassing conventional asymptotic methods.
@@
    Sample-Efficient Off-Policy Framework. We develop a sample-efficient off-policy framework that leverages the maximum entropy paradigm and importance sampling. This framework effectively manages distribution shifts and interaction data reuse, thus enhancing training stability compared to existing on-policy algorithms.
@@
    Systematic Evaluation and Robustness Analysis. We establish an evaluation platform across six benchmarks comprising stabilization and high-dimensional tracking tasks. Experiment results demonstrate superior performance, robustness and generalization of MSACL, while sensitivity analyses elucidate the multi-step bias-variance trade-off to guide diverse control tasks.
The remainder of this article is organized as follows. Section II formulates the control problems and establishes necessary theoretical foundations. Section III describes the Lyapunov certificate learning methodology. Section IV presents the optimization process within the actor-critic framework. Section V provides comprehensive experimental evaluations across six benchmarks. Finally, Section VI discusses the design philosophy and concludes this article.
@@ II Preliminaries
This section establishes the theoretical foundations for MSACL. We first define the stabilization and tracking problems for model-free discrete-time systems, then introduce Lyapunov-based exponential stability criteria, and finally review the MERL framework.
@@ II-A Problem Statement
In this article, we study the stabilization problem for general discrete-time deterministic control systems:
@@
where $\mathbf{x}_{t}\in\mathcal{X}\subseteq\mathbb{R}^{d}$ denotes the state at time $t$, $\mathbf{u}_{t}\in\mathcal{U}\subseteq\mathbb{R}^{m}$ denotes the control input, $\mathcal{X}$ and $\mathcal{U}$ represent the sets of admissible states and control inputs, respectively, and $f:\mathcal{X}\times\mathcal{U}\to\mathcal{X}$ is the flow map. Given that $f$ is unknown, we adopt a model-free control paradigm. Let $\mathcal{X}_{0}\subseteq\mathcal{X}$ denote the set of initial states. For any $\mathbf{x}_{0}\in\mathcal{X}_{0}$, let $\mathbf{x}(t,\mathbf{x}_{0})$ denote the state at time $t$ along the trajectory starting from $\mathbf{x}_{0}$. For notational simplicity, $\mathbf{x}_{t}$ is used to denote $\mathbf{x}(t,\mathbf{x}_{0})$ whenever the context is unambiguous. It is worth noting that the discrete-time case presented in this article can be similarly extended to the continuous-time case; interested readers may refer to [11] and [25] for details.
In the context of control systems, our objective is to find a state feedback policy $\pi:\mathcal{X}\to\mathcal{U}$ such that the control input $\mathbf{u}_{t}=\pi(\mathbf{x}_{t})$ renders the closed-loop system $\mathbf{x}_{t+1}=f(\mathbf{x}_{t},\pi(\mathbf{x}_{t}))$ to satisfy exponential stability. For clarity, we consider the scenario where the state $\mathbf{x}_{t}$ is fully observable, and the policy $\pi$ is stationary, i.e., it depends only on the state and not explicitly on time. Specifically, the problem studied in this article is formulated as:
@@
(Exponential Stabilization) For the control system (1), design a stationary policy $\pi$ such that for any initial state $\mathbf{x}_{0}\in\mathcal{X}_{0}$, the resulting closed-loop trajectory $\mathbf{x}(t,\mathbf{x}_{0})$ exponentially converges to the equilibrium state $\mathbf{x}_{g}$.
Without loss of generality, let $\mathbf{x}_{g}$ denote the origin, i.e., $\mathbf{x}_{g}=\mathbf{0}$. Furthermore, the tracking problem is inherently encompassed within the stabilization framework, as tracking a time-varying reference signal $\mathbf{x}_{t}^{\text{ref}}$ reduces to stabilizing the error $\mathbf{e}_{t}=\mathbf{x}_{t}-\mathbf{x}_{t}^{\text{ref}}$ at zero.
@@ II-B Exponential Stability and Lyapunov Certificates
For the control system (1), the equilibrium state $\mathbf{x}_{g}$ is said to be exponentially stable if there exist constants $C>0$ and $0<\eta<1$ such that
$\|\mathbf{x}_{t}-\mathbf{x}_{g}\|\leq C\eta^{t}\|\mathbf{x}_{0}-\mathbf{x}_{g}\|,\quad\forall\mathbf{x}_{0}\in\mathcal{X}_{0},\forall t\geq 0.$ (2)
To guarantee that the closed-loop system is exponentially stable with respect to $\mathbf{x}_{g}$, we introduce the following lemma:
@@
for all $t\in\mathbb{Z}_{\geq 0}$, where $\mathbf{x}_{t+1}=f(\mathbf{x}_{t},\mathbf{u}_{t})$, then the origin $\mathbf{x}_{g}=\mathbf{0}$ is exponentially stable. The function $V$ is referred to as an exponentially stabilizing control Lyapunov function.
Lemma 1 is a classic result in control theory, while its continuous-time counterpart is discussed in [39], we provide a concise derivation of the discrete-time version to elucidate its underlying mechanism.
Specifically, starting from the descent condition $V(\mathbf{x}_{t+1})\leq(1-\alpha_{3})V(\mathbf{x}_{t})$ in (3), the multi-step relationship can be established iteratively as:
$V(\mathbf{x}_{t+n})\leq(1-\alpha_{3})V(\mathbf{x}_{t+n-1})\leq\dots\leq(1-\alpha_{3})^{n}V(\mathbf{x}_{t}).$ (4)
By incorporating the lower and upper quadratic bounds, i.e., $\alpha_{1}\|\mathbf{x}_{t+n}\|^{2}\leq V(\mathbf{x}_{t+n})$ and $V(\mathbf{x}_{t})\leq\alpha_{2}\|\mathbf{x}_{t}\|^{2}$, the relationship between state norms across multiple steps follows as:
@@
Furthermore, by setting the initial time to $t=0$ and denoting the current time step as $n=t\in\mathbb{Z}_{\geq 0}$, we obtain the global trajectory bound:
@@
By defining the constants $C=\sqrt{\alpha_{2}/\alpha_{1}}$ and $\eta=\sqrt{1-\alpha_{3}}$, the result satisfies Definition 1, thereby proving that the control system (1) is exponentially stable with respect to $\mathbf{x}_{g}=\mathbf{0}$.